Represent the rational function f(x)=1/3x-1 using its table values and graph
1
Expert's answer
2022-04-13T13:48:39-0400

We present a table for the function "f(x)=\frac{1}{3x}-1" using "20" different values of "x". It is given below:

"\begin{array}{ll}
x&f(x)\\
-5&-1.067
\\
-4&-1.083
\\
-3&-1.111
\\
-2&-1.167
\\
-1&-1.333
\\
-0.5&-1.667
\\
-0.4&-1.833
\\
-0.3&-2.111
\\
-0.2&-2.667
\\
-0.1&-4.333
\\
0.1&2.333
\\
0.2&0.667
\\
0.3&0.111
\\
0.4&-0.167
\\
0.5&-0.333
\\1&-0.667
\\2&-0.833
\\3&-0.889
\\4&-0.917
\\
5&-0.933
\\

\end{array}"

The values for "f(x)" are rounded up to 3 decimal places. The graph of the function is presented below:



It has a vertical asymptote "x=0" . The horizontal asymptote is: "x=-1". It is due to the fact that "\frac{1}{3x}\rightarrow0,x\rightarrow\infty". The function is decaying. It is visible from the derivative: "f'=-\frac{1}{3x^2}".